OpenStudy (anonymous):

2x^12y^4z^10 divdided by 10x^9y^15z^2 dividing powers

OpenStudy (anonymous):

Divide the constants and subtract powers from smaller to bigger.$$\frac{2x^{12}y^4z^{10}}{10x^9y^{15}z^2}=\frac2{10}\times\frac{x^{12}}{x^9}\times\frac{y^4}{y^{15}}\times\frac{z^{10}}{z^2}=\frac15\times\frac{x^3}1\times\frac1{y^{11}}\times\frac{z^8}1$$Put it all together to get: \(\dfrac{x^3z^8}{5y^{11}}\)

OpenStudy (anonymous):

thank you so much!!!!!!!!!!! :)